Two 2.5 kg bowling balls are 0.50 m apart. What is the force of gravity of the first bowling ball due to the second? Compare this to the force of gravity on the second bowling ball due to the first.
1
Expert's answer
2020-12-14T07:29:39-0500

"F_{12}=F_{21}=G\frac{m_1m_2}{r^2}="


"=6.67\cdot10^{-11}\cdot\frac{2.5\cdot2.5}{0.5^2}=1.7\cdot10^{-9}(N)=1.7(nN)"